A 0.20-,uF capacitor is connected across the terminals of a variable-frequency oscillator. At what frequency is its reactance 84 Ω? What will the reactance be if the frequency is then tripled?


1
Expert's answer
2020-12-21T11:02:13-0500

The reactance of the capacitor is given by


"X_C = \dfrac{1}{2\pi f C}"

where "f" is the frequency, and "C = 0.2\mu F = 0.2\times 10^{-6}F". Since "X_C = 84\Omega", the frequency is


"f = \dfrac{1}{2\pi X_C C}\\
f = \dfrac{1}{2\pi \cdot 84 \cdot 0.2\times 10^{-6}} \approx 9.47\times 10^3Hz = 9.47\space kHz"

Answer. 9.47 kHz.